# Learning to Drive Ethically: Embedding Moral Reasoning into Autonomous Driving

### v1

@@ A.1 Hyperparameter
$R_{\text{success}}$ Success reward 25
$R_{\text{out}}$ Out of road penalty -15
$R_{\text{collision}}$1 Collision penalty -10 or -15
  • 1
    A collision with another vehicle incurs a penalty of -10, while a collision with a pedestrian or cyclist results in a penalty of -15.
A.2 Scenario analyses
@@
In this scenario, the ego vehicle has right-of-way on the main road, and the right-turning red vehicle must yield. Three baseline agents, standard mode, SACLAG, and EthicAR w/o PER, maintain a relatively high speed through the intersection. After T=1s, their TTC with the red vehicle decreases, driving up the collision risk. By T=6s, both standard mode and EthicAR w/o PER collide with the red vehicle; SACLAG narrowly avoids impact but still operates in a high-risk regime. In contrast, the EthicAR agent behaves more cautiously. At T=1s, it drifts slightly left to create space for the cyclist and settles into a comfortable speed. As the red vehicle initiates its turn, the TTC remains relatively constant, and the risks remain low. Even though the ego vehicle legally had right-of-way, EthicAR proactively yields to the red vehicle, keeping overall risk at a low level.

### v2

@@ A.1 Hyperparameter
A.2 Solutions for the longitudinal and lateral polynomial equations
As discussed in Section 4.3, for the trajectory $\xi_{e,t}$ at time step $t$, we employ a quartic polynomial in the longitudinal direction and a quintic polynomial in the lateral direction. To solve these polynomials, we use the known initial vehicle state
$l_{0},\dot{l}_{0},\ddot{l}_{0},d_{0},\dot{d}_{0},\ddot{d}_{0}]$
and the terminal state at the planning horizon $T$, given by
$\dot{l}_{T},\ddot{l}_{T},d_{T},\dot{d}_{T},\ddot{d}_{T}].$
The outputs from the RL agent provide $\dot{l}_{T}$, $d_{T}$, and $T$, while we assume $\ddot{l}_{T}=0$, $\dot{d}_{T}=0$, and $\ddot{d}_{T}=0$. Under these assumptions, the resulting trajectory minimizes jerk, thereby enhancing passenger comfort [Werling et al., 2010]. Therefore, the polynomial coefficients can be solved as follows.
Longitudinal Quartic Polynomial
The longitudinal trajectory is defined as
$l_{e}(\tau)=a_{0}+a_{1}\tau+a_{2}\tau^{2}+a_{3}\tau^{3}+a_{4}\tau^{4}.$
Using the initial state, the first three coefficients are determined:
$a_{0}=l_{0},\quad a_{1}=\dot{l}_{0},\quad a_{2}=\frac{1}{2}\ddot{l}_{0}.$
Substituting $\tau=T$ into the polynomial and its derivatives gives the terminal constraints:
$\dot{l}_{T}$ $=\dot{l}_{0}+\ddot{l}_{0}T+3a_{3}T^{2}+4a_{4}T^{3},$
$\ddot{l}_{T}$ $=\ddot{l}_{0}+6a_{3}T+12a_{4}T^{2}.$
These can be expressed in matrix form to solve for the unknown coefficients $a_{3}$ and $a_{4}$:
$\begin{bmatrix}a_{3}\\[2.0pt] a_{4}\end{bmatrix}=\begin{bmatrix}3T^{2}&4T^{3}\\[1.0pt] 6T&12T^{2}\end{bmatrix}^{-1}\begin{bmatrix}\dot{l}_{T}-\dot{l}_{0}-\ddot{l}_{0}T\\[1.0pt] \ddot{l}_{T}-\ddot{l}_{0}\end{bmatrix}$.
This yields $a_{3},a_{4}]$, completing the quartic longitudinal trajectory.
Lateral Quintic Polynomial
The lateral trajectory is defined as
$d_{e}(\tau)=b_{0}+b_{1}\tau+b_{2}\tau^{2}+b_{3}\tau^{3}+b_{4}\tau^{4}+b_{5}\tau^{5}.$
The first three coefficients are determined from the initial state:
$b_{0}=d_{0},\quad b_{1}=\dot{d}_{0},\quad b_{2}=\frac{1}{2}\ddot{d}_{0}.$
The terminal state constraints at $\tau=T$ are:
$d_{T}$ $=b_{0}+b_{1}T+b_{2}T^{2}+b_{3}T^{3}+b_{4}T^{4}+b_{5}T^{5},$
$\dot{d}_{T}$ $=b_{1}+2b_{2}T+3b_{3}T^{2}+4b_{4}T^{3}+5b_{5}T^{4},$
$\ddot{d}_{T}$ $=2b_{2}+6b_{3}T+12b_{4}T^{2}+20b_{5}T^{3}.$
These can be written in matrix form to solve for $b_{3},b_{4},b_{5}]$:
$\begin{bmatrix}b_{3}\\[1.0pt] b_{4}\\[1.0pt] b_{5}\end{bmatrix}=\begin{bmatrix}T^{3}&T^{4}&T^{5}\\[1.0pt] 3T^{2}&4T^{3}&5T^{4}\\[1.0pt] 6T&12T^{2}&20T^{3}\end{bmatrix}^{-1}\begin{bmatrix}d_{T}-b_{0}-b_{1}T-b_{2}T^{2}\\[1.0pt] \dot{d}_{T}-b_{1}-2b_{2}T\\[1.0pt] \ddot{d}_{T}-2b_{2}\end{bmatrix}$.
This provides the unknown coefficients $b_{3},b_{4},b_{5}]$, completing the lateral quintic trajectory.
A.3 Scenario analyses
@@
In this scenario, the ego vehicle has right-of-way on the main road, and the right-turning red vehicle must yield. Three baseline agents, standard mode, SACLAG, and EthicAR w/o PER, maintain a relatively high speed through the intersection. After $T=1\mathrm{s}$, their TTC with the red vehicle decreases, driving up the collision risk. By $T=6\mathrm{s}$, both standard mode and EthicAR w/o PER collide with the red vehicle; SACLAG narrowly avoids impact but still operates in a high-risk regime. In contrast, the EthicAR agent behaves more cautiously. At $T=1\mathrm{s}$, it drifts slightly left to create space for the cyclist and settles into a comfortable speed. As the red vehicle initiates its turn, the TTC remains relatively constant, and the risks remain low. Even though the ego vehicle legally had right-of-way, EthicAR proactively yields to the red vehicle, keeping overall risk at a low level.